Please help me solve this:
Let g(x)= -2(xsquared)+x+2. Find g(2) and g(1/2).
Thanks

Answer by lordacro(12)

Alright. When they say g(x) is something, and then give you g(#), they want you to fill in that number for x and find the answer.
For example,
if you have g(x)= -2(xsquared)+ x + 2, and they ask for g(2), fill in 2 everywhere you see x.
g(2)= -2(2*2) + 2 + 2
g(2)= -2(4) + 4
g(2)= -8 + 4
g(2)= -4
Try the other one on your own using what is above.